ina scientific experiment,there are 8 guinea pigs,5 of which are pregnant. if 3 are selected at random without replacement,what is the probability that all are pregnant
Answer by stanbon(75887)

in a scientific experiment, there are 8 guinea pigs,5 of which are pregnant. if 3 are selected at random without replacement, what is the probability that all are pregnant?
========
P(all pregnant) = 5C3/8C3 = (5*4*3)/(8*7*6) = 5/(2*7*2) = 5/28
-------------------
Cheers,
Stan H.